OpenStudy (anonymous):

Solve for "x". 2*3^x =7*5^x.

OpenStudy (anonymous):

1st when you typing a math problem always use the stare ( * ) when multiplying because in algebra x is a variable.

OpenStudy (anonymous):

star*

OpenStudy (anonymous):

sorry its late

OpenStudy (anonymous):

oh sorry thanks

OpenStudy (anonymous):

yeah so after you retype the question ill help

OpenStudy (anonymous):

show me first two step and im good dude.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well the first step is to get x by itself on one side

OpenStudy (jhannybean):

Is your problem : \(2\cdot 3^x=7\cdot 5^x\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1st devide 7 by 2

OpenStudy (anonymous):

The Answer is x= -2.45

OpenStudy (jhannybean):

This is a little complicated, but isolate 1 of the 2 functions with an x

OpenStudy (anonymous):

divide* jeez

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Could you just type up entire solution in one go then actually?

OpenStudy (anonymous):

here use this http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

should work

OpenStudy (anonymous):

doesn't work

OpenStudy (jhannybean):

Actually!!!

OpenStudy (jhannybean):

Start by taking the log of both sides.

OpenStudy (anonymous):

thats weird it always does

OpenStudy (jhannybean):

\[2\cdot 3^x = 7\cdot 5^x\]Taking the log of both sides..\[\log(2\cdot 3^x) = \log(7\cdot 5^x)\]With the properties of logs, you know that \(\log(a) +\log(b) = \log(ab)\) So apply this rule.

OpenStudy (jhannybean):

Apply this rule to both sides to separate the functions. \[\log(2) +\log(3^x) = \log(7) +\log(5^x)\]Now you can apply the power rule to log functions here, also. \[\log(x^n) = n\log(x)\]

OpenStudy (jhannybean):

Are you with me so far?

OpenStudy (anonymous):

thank you! :)